OpenStudy (anonymous):

If cot θ = −1/ 3 and cos θ < 0, compute csc θ and sin θ.

OpenStudy (anonymous):

cot= -1/3

OpenStudy (anonymous):

 csc(t) is 1/sin(t)= H/O = hyp./Opp. sin(t) is O/H or opp./hyp.

OpenStudy (anonymous):

Hyp = sqrt( (1)(1) + (-3)(-3) ) =sqrt(10)

OpenStudy (anonymous):

So what is cdc theta and sin theta?

OpenStudy (anonymous):

csc*